OpenStudy (anonymous):

*Fan & medal* >ONE QUESTION< Ball 1, with a mass of 10 kg, is moving to the right at 20 m/s. Ball 2 is moving to the left at 20 m/s. Upon collision, Ball 2 comes to a standstill, and Ball 1 moves to the left at twice its original speed. What is the mass of Ball 2?

OpenStudy (anonymous):

@Dbzfan836 This is "perfectly elastic collision" right?

OpenStudy (dbzfan836):

yes.

OpenStudy (anonymous):

Ok, ima try to figure it out.

OpenStudy (anonymous):

Yeah I have no idea how to find the mass. e.e

OpenStudy (anonymous):

Wait is the mass 30 by chance?

OpenStudy (dbzfan836):

no.

OpenStudy (dbzfan836):

You find mass by taking the momentum of the object, and dividing that by its velocity

OpenStudy (anonymous):

Its momentum is 0 right?

OpenStudy (anonymous):

Nvm, 0 is its end velocity.

OpenStudy (dbzfan836):

Alright, to find its momentum, you have to do the same kind of calculations from last time.

OpenStudy (anonymous):

Ok so I do 10 X 20 = 200 right?

OpenStudy (anonymous):

Don't I need its mass to do the same calculations as last time?

OpenStudy (dbzfan836):

yes, so do all the other object's masses

OpenStudy (dbzfan836):

in the end, what was ball 1's mass?

OpenStudy (dbzfan836):

I mean momentum

OpenStudy (dbzfan836):

in the end, what was ball 1's momentum?

OpenStudy (anonymous):

As in any collision momentum is always conserved Use p total before = p total after p total after can be found using p = mv (since only one ball is in motion we can ignore the second ball) p= 10 * 40 =400 kg* m/s p total after must also be 400 p total before = p ball1 + p ball2 m1v1 + m2v2 isolate for the mass of ball 2 m2= (p total - m1v1) / v2 = (400 - 10(20))/ 20 = 200/20 = 10 kg

OpenStudy (dbzfan836):

dude, you have to let the guy learn from this. You have to let him try

OpenStudy (anonymous):

I showed him the work... Was I wrong to do that? Im sorry...:(

OpenStudy (anonymous):

Y'all both are helping, ether way is fine. Its just better if someone explains it to me lol.

OpenStudy (anonymous):

Okay. My bad....:/

OpenStudy (dbzfan836):

@JahnkeSwim I mean it is wrong to just give answeres

OpenStudy (anonymous):

The mass is not 10 though, I have a lab simulation I have to enter the answer in. It says 10 is not right.

OpenStudy (dbzfan836):

alright, did you put 10 Kg?

OpenStudy (dbzfan836):

or just 10?

OpenStudy (anonymous):

It automatic puts the kg on it. I have to find the mass of Ball 2.

OpenStudy (dbzfan836):

well, that means we have to start over this, writing it all down to make sure where we were wrong.

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

So this is what we know. Ball 1 Mass: 10 kg Velocity: 20 m/s Ball 2 Mass: unkown Velocity: 20 m/s

OpenStudy (dbzfan836):

yes

OpenStudy (anonymous):

Ok so we take 10kg and 20 m/s and multiply it together to get 200 right?

OpenStudy (dbzfan836):

yes

OpenStudy (anonymous):

Ok so what next? We don't know the mass of Ball 2 so idk what to do... o.o

OpenStudy (dbzfan836):

alright, what is the mass after the collision?

OpenStudy (anonymous):

Idk...

OpenStudy (anonymous):

I put in 30 and it says its right, I'm just gonna put 30 as the answer even though I can't show my work. x_x

OpenStudy (anonymous):

Thanks for all the help. :)